Director's Time

Observers: B.Rothberg, T. Rochais (mountain), O. Kuhn, M. (Tucson)

ISA: B.Rothberg
Telescope Operator: D.Gonzalez-Huerta (LBTO)
AO Support: D.Miller
Telescope Support: M.Wagner (LBTO)
Software Support: S.Hooper
Instrument Support: J.Morris (LBTO)
Engineering Support: D.Ashby

Summary

First Half D-time

Details

UT 01:05 -Starting up LUCI - taking 5 4sec test darks

UT 01:07 - Running Dark Script for O.Kuhn

UT 01:23 - Strange behavior with NX Client and LUCI - the CTRL key doesn't function - same issue with logging in as engineer

UT 06:12 - behavior with NX due to issue with mouse control setting in Gnome Desktop - choice to have CTRL key show mouse position was selected

UT 06:13 - Still closed, Humidity, Dew Point, Wind in Yellow.

UT 07:18 - End of D-time - No data taken

-- %USERSIG{BarryRothberg - 2015-05-23}%

Comments

 
Topic revision: r2 - 23 May 2015, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback